OpenStudy (anonymous):

4/x+3 - x/x-3 - 18/x^2-9

OpenStudy (anonymous):

add or subtract

OpenStudy (anonymous):

hint the common denominator for \(x+3,x-3,x^2-9\) is \(x^2-9\)

OpenStudy (anonymous):

so i would mutiply x62-9 to all the number on the top

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the first you would multiply top and bottom by \(x-3\) the second by \(x+3\) and the third you would leave alone since it is already over \(x^2-9\)

OpenStudy (anonymous):

4x+12/x2-9 - x^2-3x/x^2-9-18/x^2-9

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

you got them backwards

OpenStudy (anonymous):

\[\frac{4}{x+3}=\frac{4(x-3)}{(x+3)(x-3)}\]

OpenStudy (anonymous):

\[\frac{x}{x-3}=\frac{x(x+3)}{(x+3)(x-3)}\]

OpenStudy (anonymous):

would i cross out the (x-3)

OpenStudy (anonymous):

oh no!!

OpenStudy (anonymous):

+ i mean

OpenStudy (anonymous):

you are builing up the fractions so they have the same denominator, just like with numbers

OpenStudy (anonymous):

*building

OpenStudy (anonymous):

yeah im stuck

OpenStudy (anonymous):

-x2−x+30/(x+3)(x−3) my friend got this but i dont know how he did it

OpenStudy (anonymous):

@sweetburger

OpenStudy (sweetburger):

factor out a negative in the numerator then proceed to factor the rest of the numerator

OpenStudy (sweetburger):

actually if thats the last step u have i dont think anything else cancels you might want someone to confirm

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well if you go back and read what has been posted... about using common denominators the problem is \[\frac{4(x -3)}{(x -3)(x+3)} - \frac{x(x +3)}{(x -3)(x+3) }- \frac{18}{(x -3)(x+3)}\] now just distribute and collect like terms in the numerator the see what you have

OpenStudy (anonymous):

so i dont cross out anything